dx/dt=x-4y



dy/dt=x+y

1
Expert's answer
2021-11-29T11:36:42-0500

"Let\\
\frac{d}{dt}=D.\\
\text{The system of equation is given by,}\\
Dx=x-4y\\
\implies(D-1)x+4y=0\space ------(1)\\
And\\
Dy=x+y\\
\implies -x+(D-1)y=0\space ------(2)\\
\text{First Multiply (2) with (D-1) and add(1),we get}\\
[4+(D-1)^2]y=0\\
(D^2-2D+5)y=0
\text{Auxiliary equation,}\\
m^2-2m+5=0\\
m=\frac{2\pm \sqrt{-16}}{2}\\
=1\pm 2i\\
\text{The solution is given by,}\\
y(t)=e^t(Acos(2t)+Bsin(2t))\space ------(3)\\
\text{Putting the value of y in (2), we get}\\
-x+(D-1)(e^t(Acos(2t)+Bsin(2t))=0\\
x=e^t(-2Asin(2t)+2Bcon(2t))+e^t(Acos(2t)+Bsin(2t))-e^t(Acos(2t)+Bsin(2t))\\
x(t)=e^t(-2Asin(2t)+2Bcon(2t))\space ------(4)\\
\text{Thus, (3) and (4) are the required solution.}"